OpenStudy (anonymous):

Differentiate the function y = f(x) = (x^2−x)/x

OpenStudy (anonymous):

Do you know the quotient rule?

OpenStudy (anonymous):

No. I think that may be what I need to do a few of the derivative problems I can't understand

OpenStudy (anonymous):

If the function is rational: \[\frac{g(x)}{h(x)}\] Then the differentiation is: \[\frac{g'(x)h(x) - g(x)h'(x)}{[h(x)]^2}\]

OpenStudy (anonymous):

So here you have: \[\frac{(x^2 - x)' x - (x^2 - x)x'}{x^2}\]

OpenStudy (anonymous):

Using that, I get x^2/x^2. Can 1 be my answer?

OpenStudy (anonymous):

Yep. That's it.

OpenStudy (tamtoan):

if you don't have to use quotient rule, try this : f(x) = (x^2 - x)/x = x^2/x - x/x = x -1 so f'(x) = 1

OpenStudy (anonymous):

Thanks!